OpenStudy (anonymous):

20x^6-4x^4+8 / -4x Divide!!!!!

OpenStudy (anonymous):

-5x^5.. etc

OpenStudy (anonymous):

then I was way off lol cause I got 13x^5+-8x^3+4x as the answer

OpenStudy (anonymous):

You should be able to reverse it ie multiply your answer by what you divided and get what you started with.

OpenStudy (anonymous):

is it -5x^5+1^3-2x?

OpenStudy (anonymous):

Well, that doesn't reverse, does it?

OpenStudy (zepp):

http://puu.sh/s9dp :D

OpenStudy (anonymous):

no, man I am horrible at math......

OpenStudy (zepp):

Since \[\frac{a+b+c}{d}\]could also be written as \[\frac{a}{d}+\frac{b}{d}+\frac{c}{d}\]

OpenStudy (anonymous):

Excuses, excuses...Zepp will show you.

OpenStudy (zepp):

lol, don't be so rough, @estudier :P

OpenStudy (anonymous):

-5c^5+1x^3 is that right so far?

OpenStudy (zepp):

Yes.

OpenStudy (zepp):

Where does the c come from? D:

OpenStudy (anonymous):

-5x^5+1x^3-2/x

OpenStudy (anonymous):

c idk where it comes from.

OpenStudy (zepp):

Yep, that's right :) See, you aren't so bad at math, stop underestimating yourself ;D

OpenStudy (anonymous):

\[ 20x^6-4x^4+8\over-4x\\ {20x^6\over-4x}+{-4x^4\over-4x}+{8\over-4x}\\ -5x^5+x^3+{-2\over x} \] If that helps.

OpenStudy (anonymous):

thanks Zepp